Decreasing the time constant of a RC filter increases the cut-off frequency.

True

False
Answer #1

time constant of RC filter = R×C

And cut - off frequecy of RC filter = 1/(2×π×R×C)

or f​​​​​​c​​​( cut off frequency ) = 1/(2×π×time constant)

hence if time constant decreases then cut off frequency increases.

this statement is TRUE .